# Normalizing Flows are Capable Generative Models

### v1

@@ 2.6 Guidance
Unconditional Guidance
In addition, we also introduce a novel method for guiding unconditional models. The basic idea is to construct predictions of inferior quality, analogous to the role of unconditional ones. In order to do so, we override the notation yet again and introduce $\mu_{i}^{t}(\cdot;\tau),\alpha_{i}^{t}(\cdot;\tau)$. Here we have let the predictions $\mu_{i}^{t},\alpha_{i}^{t}$ take an additional parameter $\tau$, which corresponds to a manually injected temperature term to all the attention layers in the Transformer for $f^{t}$. Namely, for each attention layer in $f^{t}$, we divide the attention logits by $\tau$, before normalizing it with the Softmax function. A $\tau$ larger or smaller than 1 makes the attention overly smooth or sharp, either way reducing the Transformer’s ability to correctly predict the next variable’s transformations.
@@ 3 Experiments
Our model significantly advances the state-of-the-art in likelihood modeling, achieving substantial improvements over previous methods. At the same time, TarFlow demonstrates unprecedented sample generation quality for the class of normalizing flow models, reaching levels comparable to diffusion models and GANs for the first time. Detailed experimental configurations are provided in the Appendix.
@@ 3.1 Likelihood
Model Type FID $\downarrow$
MFM  (Pooladian et al., 2023) Diff/FM 11.82
FM  (Lipman et al., 2023b) Diff/FM 13.93
AGM  (Chen et al., 2024) Diff/FM 10.07
IC-GAN (Casanova et al., 2021) GAN 10.40
Self-sup GAN (Noroozi, 2020) GAN 19.20
TarFlow [2-768-8-8-$\mathcal{N}(0,0.05^{2})$] (Ours) NF 18.42
@@
Figure 4: Left: The effect of input noise $\sigma$ and denoising, all samples are generated with guidance weight $w=2$ on ImageNet 128x128 from the same initial noise, better viewed when zoomed in. Right: Sample FID vs input noise $\sigma$ on ImageNet 64x64, with and without denoising. Before denosing, it first appears that small $\sigma$ has the best FID, due to the smaller amount of noise present in the raw samples. However, after denoising with Equation 8, slightly larger $\sigma$ favors better FID and demonstrates more consistent shapes. Note that the scale of the right y-axis differs from that of the left.
@@
3.2 Generation
We then perform quantitative evaluations in terms of FID, on the ImageNet models. For each setting, we randomly generate 50K samples, and compare it with the statistics from the entire training set. We search for the best guidance weights (and attention temperature in the unconditional case). The results are summarized in Table 3, 4, 5. In all settings, we see that TarFlow produces competitive FID numbers, often times better than strong GAN baselines, and approaching results from recent Diffusion Models. It is also interesting to note that we found no publicly reported NF based FID numbers on the ImageNet level datasets, most likely due to the lack of presentable results from the NF community.
3.3 Ablation on Noise Augmentation and Denoising
Next, we experiment with different Gaussian noise levels $\sigma$ during training on class conditional ImageNet 64x64. We use an architecture configuration of 4-1024-8-8, and vary $\sigma$ in {0.01, 0.05, 0.2, 0.5}. For fast experimentation, we train all models for only 100 epochs with a batch size of 512. We evaluate these models with a guidance $w=2$ and plot the 50K sample FIDs before and after the score based denoising. For visual comparison, we also train two models ImageNet 128x128 models, with the architecture 4-1024-8-8 and noise $\sigma\in[0.05,0.15]$, respectively. We show the FID curves together with the visual examples in Figure 4. There are two important observations. First, naively increasing the noise level on the surface appears to hurt the raw samples’ quality. However, this is no longer the case after applying the denoising step in Equation 8. Denoising successfully cleans up the noisy raw samples, and as a result the best visual quality occurs at a moderate (but still relatively small) amount of noise. This verifies the necessity of our proposed sampling procedure, whereas the combination of noise augmented training and the score based denoising step work organically together to produce the best generative capability.
3.4 Ablation on Guidance
(a) Refer to caption
(b) Refer to caption
Figure 7: (a) A typical training run on ImageNet 64x64. Our loss smoothly decreases during training, and is positively correlated with FID. (b) Depth configuration (in the form of $T\times K$) vs training loss. Overall, we see a strong positive correlation between the training loss and FID. Left, the optimal training loss happens when the capacity is evenly allocated to number of blocks and number of layers per block. Interestingly, the special case of 1 block degenerates to an incapable model, which has both high loss and FID equivalent to random guess. Right, increasing both the number of blocks and number of layers per block improves the model’s loss and sampling quality.
We then turn our eyes to guidance. Similar to CFG in Diffusion Models, guidance for TarFlow is a post training technique that allows us to vary the model’s sample quality during inference. We perform qualitative and quantitative evaluations on both the class conditional and unconditional versions of ImageNet 64x64. The results are shown in Figure 5 and 6. In terms of FID, the guidance weight $w$ plays an effective role for both models. Visually, it is also clear that guidance allows the model to converge to more recognizable modes, presenting more aesthetic samples. Interestingly, this is also somewhat true for the unconditional models, whereas both the guidance weight $w$ and attention temperature $\tau$ contribute to the degree of guidance. We show more guidance comparisons in the Appendix.
3.5 Ablation on Model Scaling
Another important aspect of our work is the model’s scaling properties. We first show a typical training loss curve together with an online monitoring of the model’s sample quality in terms of FID (we use 4096 samples for efficiency). This is shown in Figure 7(a). We see that the loss curve is smooth and monotonic, and it has a strong positive correlation with the FID curve.
@@ 3.5 Ablation on Model Scaling
3.6 Visualizing Sample Trajectory
Lastly, thanks to the residual style composition of TarFlow, we can also visualize the generation process by reshaping each $\{z^{t}\}$ to the pixel space. We visualize two sampling sequences with the ImageNet 128x128 model in Figure 8. Interestingly, the sample trajectories highly resemble those from a Diffusion model, in the sense that the initial noise is gradually transformed into visible inputs – though they are trained with completely different objectives.
4 Related Work
Coupling-based Normalizing Flows
Many research endeavors focus on developing expressive invertible transformations with tractable Jacobian computations for constructing normalizing flows. NICE (Dinh et al., 2014) introduced additive coupling layers to construct the transformations and simplified the computation of the Jacobian determinant. RealNVP (Dinh et al., 2017) extended this approach by incorporating scaling and shifting operations to enhance the model’s expressiveness. Glow (Kingma & Dhariwal, 2018) advanced these models by introducing invertible $1\times 1$ convolutions, achieving improved results in image generation tasks. Flow++ (Ho et al., 2019) further introduced attention mechanisms to enhance the model’s expressiveness. iResNet (Behrmann et al., 2019) demonstrated the invertibility of standard ResNet architectures (He et al., 2016) by incorporating a normalization step. Moreover, normalizing flows have played an important role in improving the performance of Variational Autoencoders (VAEs) (Kingma & Welling, 2014) by providing flexible posterior distributions (Su & Wu, 2018; Zhang et al., 2020), and in diffusion models (Song et al., 2021b) by introducing adaptable nonlinear forward and backward diffusion and drift terms (Kim et al., 2022). What’s shared in common among these designs is that they need carefully wired and restrictive architectures, which poses great a challenge in scaling the model’s capacity.
Continuous Normalizing Flows
Neural Ordinary Differential Equations (Chen et al., 2018) reformulate the standard ResNet architecture as a deterministic ordinary differential equation in the continuous-time limit. This formulation can be naturally extended to normalizing flows, resulting in Continuous Normalizing Flows. In this framework, the invertibility of the network is inherently satisfied, and the computation of the Jacobian determinant within the normalizing flow is reduced to calculating the trace of the Jacobian. By leveraging Pontryagin’s Maximum Principle (Pontryagin, 2018), gradient backpropagation through such networks can be efficiently performed using the adjoint method, which requires only $\mathcal{O}(1)$ memory complexity. FFJORD (Grathwohl et al., 2019) further simplifies the expensive Jacobian computation by employing Hutchinson’s trace estimator (Hutchinson, 1989). However, these models often suffer from numerical instability during training and sampling, which has been extensively analyzed in (Zhuang et al., 2021; Liu et al., 2021). The expressive capability can be further improved by augmenting auxiliary variables (Dupont et al., 2019; Chalvidal et al., 2021). In comparison, TarFlow enables an unconstrained architecture design paradigm by fully taking advantage of the power of causal Transformers, which we believe is a key component for realizing the true potential of the NF principle.
Autoregressive Normalizing Flows
There have also been significant efforts to integrate normalizing flows with autoregressive models. IAF (Kingma et al., 2016) introduced dimension-wise affine transformations conditioned on preceding dimensions for variational inference, and MAF (Papamakarios et al., 2017) leveraged the MADE (Germain et al., 2015) architecture to construct invertible mappings through autoregressive transformations. Neural autoregressive flow (Huang et al., 2018) replaces the affine transformation in MAF by parameterizing a monotonic neural network for each dimension to enhance the transformation’s expressiveness, at the cost of losing analytical invertibility. T-NAF (Patacchiola et al., 2024) extends NAF by introducing a single autoregressive Transformer. Block Neural Autoregressive Flow (Cao et al., 2019) takes a different approach by fitting an end-to-end autoregressive monotonic neural network, rather than NAF’s dimension-wise sequence parameterization, but also sacrifices analytical invertibility. TarFlow differs from these as we show that it is sufficient to stack multiple iterations of block autoregressive flows with standard Transformer model in alternating directions, without the need for other types of flow operations.
Probability Flow in Diffusion
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) generate data by simulating Stochastic Differential Equations. Song et al. (2021b) provided a deterministic Ordinary Differential Equation counterpart to this generative approach, also known as the scoreflow (Song et al., 2021a). The scoreflow can be understood as a special case of continuous normalizing flows by interpreting the learned score term and base drift as a new parameterized drift term. However, Lu et al. (2022) demonstrated that the training objective in diffusion models, which relies on first-order score approximation, is not sufficient to maximize the likelihood of the scoreflow. TarFlow differs from these instances as it is directly trained with the MLE objective, without the need for excessively large Gaussian noise during training.
Autoregressive Models for Image Generation
Diffusion models, other generative models
Diffusion models (Ho et al., 2020; Song et al., 2021b) are emerging generative models that achieve appealing results. Stable Diffusion (Podell et al., 2024) and OpenSora (Zheng et al., 2024) push the boundaries of diffusion models’ capabilities, demonstrating their ability to generate extremely high-dimensional data. Besides, Variational Autoencoders (VAEs) (Kingma & Welling, 2014) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) are also popular generative models. By avoiding the posterior collapse issue, VQ-VAE (van den Oord et al., 2017) demonstrates impressive generative performance and subsequently serves as an essential component in the later latent diffusion model (Podell et al., 2024). In the realm of GANs, Karras et al. (2019); Kang et al. (2023); Brock et al. (2019) showcase the remarkable capability of GANs to generate high-resolution images with comparatively cheap inference costs, though the training stability of GANs remains challenging (Wiatrak et al., 2019). TarFlow represents an orthogonal learning paradigm to these methods, with its unique benefits and challenges.
5 Conclusion
6 Acknowledgements
Appendix A Experimental details
Table 6: Hyper parameters for the best performing model on each task.
Cond ImageNet 64x64 2 768 8 8 $\mathcal{N}(0,0.05^{2})$ 256 200 8
Cond ImageNet 128x128 4 1024 8 8 $\mathcal{N}(0,0.15^{2})$ 768 320 32
Cond AFHQ 256x256 8 768 8 8 $\mathcal{N}(0,0.07^{2})$ 256 4000 8
Appendix B Inference Implementation
Another component in our sampling pipeline is the score based denoising step. The time of this step is equal to two forward model calls, which usually happens in a matter of seconds. A practical bottlenck is that this step is more memory consuming than the flow reverse step, due to the need of caching all intermediate activations for back propagation. In theory, this can be further alleviated by adopting techniques like gradient checkpointing, essentially trading time for memory.
Appendix C Additional samples

### v3

@@ 2.5 Score Based Denoising
It is worth noting that our score based denoising technique uses only the TarFlow model itself, without requiring any extra modules. The fact that this works well (as demonstrated later in Sec. 3.3) suggests that learning the density of the noisy distribution is sufficient for recovering the score function, which is an interesting and significant result on its own in the context of score based generative models.
@@ 2.6 Guidance
We have also discovered as method for applying guidance to unconditional models, see Appendix Section B for details.
@@ 3.1 Likelihood
TarFlow [2-1024-8-8-$\mathcal{N}(0,0.05^{2})$] (Ours) NF 2.66
@@
Figure 4: Top: The effect of input noise $\sigma$ and denoising, all samples are generated with guidance weight $w=2$ on ImageNet 128x128 from the same initial noise, better viewed when zoomed in. Bottom: Sample FID vs input noise $\sigma$ on ImageNet 64x64, with and without denoising. Before denosing, it first appears that small $\sigma$ has the best FID, due to the smaller amount of noise present in the raw samples. However, after denoising with Equation 8, slightly larger $\sigma$ favors better FID and demonstrates more consistent shapes. Note that the scale of the right y-axis differs from that of the left.
@@
3.2 Generation
We then perform quantitative evaluations in terms of FID, on the ImageNet models. For each setting, we randomly generate 50K samples, and compare it with the statistics from the entire training set. We search for the best guidance weights (and attention temperature in the unconditional case). The results are summarized in Table 3, 6, 4. In all settings, we see that TarFlow produces competitive FID numbers, often times better than strong GAN baselines, and approaching results from recent Diffusion Models. It is also interesting to note that we found no publicly reported NF based FID numbers on the ImageNet level datasets, most likely due to the lack of presentable results from the NF community.
(a) Refer to caption
(b) Refer to caption
Figure 6: (a) A typical training run on ImageNet 64x64. Our loss smoothly decreases during training, and is positively correlated with FID. (b) Depth configuration (in the form of $T\times K$) vs training loss. Overall, we see a strong positive correlation between the training loss and FID. Left, the optimal training loss happens when the capacity is evenly allocated to number of blocks and number of layers per block. Interestingly, the special case of 1 block degenerates to an incapable model, which has both high loss and FID equivalent to random guess (FID 267). Right, increasing both the number of blocks and number of layers per block improves the model’s loss and sampling quality.
3.3 Ablation on Noise Augmentation and Denoising
Next, we experiment with different Gaussian noise levels $\sigma$ during training on class conditional ImageNet 64x64. We use an architecture configuration of 4-1024-8-8, and vary $\sigma$ in {0.01, 0.05, 0.2, 0.5}. For fast experimentation, we train all models for only 100 epochs with a batch size of 512. We evaluate these models with a guidance $w=2$ and plot the 50K sample FIDs before and after the score based denoising. For visual comparison, we also train two models ImageNet 128x128 models, with the architecture 4-1024-8-8 and noise $\sigma\in[0.05,0.15]$, respectively. We show the FID curves together with the visual examples in Figure 4. There are two important observations. First, naively increasing the noise level on the surface appears to hurt the raw samples’ quality. However, this is no longer the case after applying the denoising step in Equation 8. Denoising successfully cleans up the noisy raw samples, and as a result the best visual quality occurs at a moderate (but still relatively small) amount of noise. This verifies the necessity of our proposed sampling procedure, whereas the combination of noise augmented training and the score based denoising step work organically together to produce the best generative capability. See the Appendix for more visualizations of the effect of denoising.
3.4 Ablation on Guidance
To see the effect of guidance, we perform qualitative and quantitative evaluations on both the class conditional and unconditional versions of ImageNet 64x64. The results are shown in Figure 5 and 7 (in Appendix). In terms of FID, the guidance weight $w$ plays an effective role for both models. Visually, it is also clear that guidance allows the model to converge to more recognizable modes, presenting more aesthetic samples. Interestingly, this is also somewhat true for the unconditional models, whereas both the guidance weight $w$ and attention temperature $\tau$ contribute to the degree of guidance. We show more guidance comparisons in the Appendix.
3.5 Ablation on Model Scaling
Regarding scaling, we first show a typical training loss curve together with an online monitoring of the model’s sample quality in terms of FID (we use 4096 samples for efficiency). This is shown in Figure 6(a). We see that the loss curve is smooth and monotonic, and it has a strong positive correlation with the FID curve.
3.6 Comparison with VP and Channel Coupling
We also ablation two important design choices of TarFlow: the non-volume preserving (NVP) and autoregressive aspect. We train two baseline models: one to change NVP to VP (done by setting $\alpha^{t}_{i}(\cdot)$ to 00 in Equation 3); and the other a channel coupling baseline by using the same architecture but removing the causal masks. We use the same 0.05 Gaussian noise on both, and found that and denoising consistently help. For the VP baseline, we found that we also need to learn the prior’s variance as the latents tend to have very small magnitudes which makes sampling from the standard Gaussian prior immediately fail. All models share similar training costs which makes it a fair comparison. The results are shown in Table 5. We see that both variants significantly under perform TarFlow. Also, interestingly, guidance consistently improves both settings.
Table 5: VP and channel coupling w.r.t. FID on ImageNet 64$\times$64.
Guidance TarFlow VP channel coupling
0 25.3 81.5 50.3
2 5.7 51.0 20.4
4 Related Work
Coupling-based Normalizing Flows. NICE (Dinh et al., 2014) introduced additive coupling layers to construct the transformations and simplified the computation of the Jacobian determinant. RealNVP (Dinh et al., 2017) extended this approach by incorporating scaling and shifting operations to enhance the model’s expressiveness. Glow (Kingma & Dhariwal, 2018) advanced these models by introducing invertible $1\times 1$ convolutions, achieving improved results in image generation tasks. Flow++ (Ho et al., 2019) further introduced learned dequantization noise, a sophisticated coupling layer and attention mechanisms to enhance the model’s expressiveness. What’s shared in common among these designs is that they need carefully wired and restrictive architectures, which poses great a challenge in scaling the model’s capacity.
Continuous Normalizing Flows. Neural Ordinary Differential Equations (Chen et al., 2018) based Continuous Normalizing Flows is an alternative NF design principle. In this framework, the invertibility of the network is inherently satisfied, and the computation of the Jacobian determinant within the normalizing flow is reduced to calculating the trace of the Jacobian. FFJORD (Grathwohl et al., 2019) further simplifies the expensive Jacobian computation by employing Hutchinson’s trace estimator (Hutchinson, 1989). However, these models often suffer from numerical instability during training and sampling, which has been extensively analyzed in (Zhuang et al., 2021; Liu et al., 2021). The expressive capability can be further improved by augmenting auxiliary variables (Dupont et al., 2019; Chalvidal et al., 2021). In comparison, TarFlow enables an unconstrained architecture design paradigm by fully taking advantage of the power of causal Transformers, which we believe is a key component for realizing the true potential of the NF principle.
Autoregressive Normalizing Flows. IAF (Kingma et al., 2016) introduced dimension-wise affine transformations conditioned on preceding dimensions for variational inference, and MAF (Papamakarios et al., 2017) leveraged the MADE (Germain et al., 2015) architecture to construct invertible mappings through autoregressive transformations. Neural autoregressive flow (Huang et al., 2018) replaces the affine transformation in MAF by parameterizing a monotonic neural network, at the cost of losing analytical invertibility. T-NAF (Patacchiola et al., 2024) extends NAF by introducing a single autoregressive Transformer. Block Neural Autoregressive Flow (Cao et al., 2019) fits an end-to-end autoregressive monotonic neural network, rather than NAF’s dimension-wise sequence parameterization, but also sacrifices analytical invertibility. TarFlow differs from these as we show that it is sufficient to stack multiple iterations of block autoregressive flows with standard Transformer model in alternating directions, without the need for other types of flow operations.
Probability Flow in Diffusion & Flow Matching. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) generate data by simulating Stochastic Differential Equations. Song et al. (2021) provided a deterministic Ordinary Differential Equation (ODE) counterpart to this generative approach, also known as the probability flow (Song et al., 2021). Similarly, Flow Matching (Lipman et al., 2023a) proposes to learn such ODEs by training on linear interpolants of data and noise with the velocity prediction objective. Importantly, TarFlow differs from these instances as it is directly trained with the MLE objective, without the need for excessively large Gaussian noise during training.
See also Sec. A in Appendix for additional related works.
5 Conclusion
Acknowledgements
Impact Statement
This paper concerns the generative modeling methodology. While we do not see immediate societal implications from our technical contribution, there are potential impacts when it is used in training foundational generative models.
Appendix A Additional Related Work
Autoregressive Models for Image Generation
Diffusion models, other generative models
Diffusion models (Ho et al., 2020; Song et al., 2021) are emerging generative models that achieve appealing results. Stable Diffusion (Podell et al., 2024) and OpenSora (Zheng et al., 2024) push the boundaries of diffusion models’ capabilities, demonstrating their ability to generate extremely high-dimensional data. Besides, Variational Autoencoders (VAEs) (Kingma & Welling, 2014) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) are also popular generative models. By avoiding the posterior collapse issue, VQ-VAE (van den Oord et al., 2017) demonstrates impressive generative performance and subsequently serves as an essential component in the later latent diffusion model (Podell et al., 2024). In the realm of GANs, Karras et al. (2019); Kang et al. (2023); Brock et al. (2019) showcase the remarkable capability of GANs to generate high-resolution images with comparatively cheap inference costs, though the training stability of GANs remains challenging (Wiatrak et al., 2019). TarFlow represents an orthogonal learning paradigm to these methods, with its unique benefits and challenges.
Appendix B Guidance
In addition to conditional guidance, we also introduce a novel method for guiding unconditional models. The basic idea is to construct predictions of inferior quality, analogous to the role of unconditional ones. In order to do so, we override the notation yet again and introduce $\mu_{i}^{t}(\cdot;\tau),\alpha_{i}^{t}(\cdot;\tau)$. Here we have let the predictions $\mu_{i}^{t},\alpha_{i}^{t}$ take an additional parameter $\tau$, which corresponds to a manually injected temperature term to all the attention layers in the Transformer for $f^{t}$. Namely, for each attention layer in $f^{t}$, we divide the attention logits by $\tau$, before normalizing it with the Softmax function. A $\tau$ larger or smaller than 1 makes the attention overly smooth or sharp, either way reducing the Transformer’s ability to correctly predict the next variable’s transformations.
Model Type FID $\downarrow$
MFM  (Pooladian et al., 2023) Diff/FM 11.82
FM  (Lipman et al., 2023b) Diff/FM 13.93
AGM  (Chen et al., 2024) Diff/FM 10.07
IC-GAN (Casanova et al., 2021) GAN 10.40
Self-sup GAN (Noroozi, 2020) GAN 19.20
TarFlow [2-768-8-8-$\mathcal{N}(0,0.05^{2})$] (Ours) NF 18.42
Appendix C Experimental details
Table 7: Hyper parameters for the best performing model on each task.
Cond ImageNet 64x64 2 1024 8 8 $\mathcal{N}(0,0.05^{2})$ 768 320 32
Cond ImageNet 128x128 4 1024 8 8 $\mathcal{N}(0,0.15^{2})$ 768 320 32
Cond AFHQ 256x256 8 768 8 8 $\mathcal{N}(0,0.07^{2})$ 256 4000 8
Appendix D Inference Implementation
Another component in our sampling pipeline is the score based denoising step. The time of this step is equal to two forward model calls, which usually happens in a matter of seconds. A practical bottleneck is that this step is more memory consuming than the flow reverse step, due to the need of caching all intermediate activations for back propagation. In theory, this can be further alleviated by adopting techniques like gradient checkpointing, essentially trading time for memory.
@@ Appendix D Inference Implementation
D.1 Visualizing Sample Trajectory
Thanks to the residual style composition of TarFlow, we can also visualize the generation process by reshaping each $\{z^{t}\}$ to the pixel space. We visualize two sampling sequences with the ImageNet 128x128 model in Figure 8. Interestingly, the sample trajectories highly resemble those from a Diffusion model, in the sense that the initial noise is gradually transformed into visible inputs – though they are trained with completely different objectives.
@@ D.1 Visualizing Sample Trajectory
Appendix E Additional samples